# A decoder-only foundation model for time-series forecasting

### v2

@@ Appendix A Appendix
A.1 Metrics
@@
A.2 Additional Empirical Results
In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1
A.2.1 Darts
We present the MAE results individually from all 8 datasets in Table 1. It can be seen that TimesFM performs well for all datasets with clear seasonal patterns. We do not perform the among the top-3 only in Sunspot and Wooly datasets. On an average we are within significant level of the best model and a close second on the mean scaled MAE metric.
@@ A.2.1 Darts
gp arima TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) NAIVE
AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 14.75 81.45
AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 10.25 96.35
GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.69 2.29
MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 22.46 85.71
SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 50.88 48.24
WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2462.11 4075.28
WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 917.10 1210.33
HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.44 5.92
A.2.2 Monash
@@
llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
@@
australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 523.12 659.60
bitcoin 1.75e18 5.33e18 5.33e18 9.9e18 1.10e18 3.62e18 6.66e18 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.97e18 5.32e18
pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 58.35 170.88
weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.17 2.36
nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 4.09 8.26
nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 14.33 16.71
tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 93572.42 99456.05
tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 10037.84 15845.10
tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 2916.02 5636.83
cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 896414.41 386526.37
covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 555.24 2653.98
fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 1823.16 2825.67
traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.03
traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.13 1.19
saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 24.53 21.50
us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 408.49 1152.67
hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 19.34 24.07
solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1184.01 1729.41
Table 3: We present the mean msMAPE results for our methods along size monash baselines. We also include the naive baseline that predicts the last values in the context repeatedly.
llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer TimesFM(ZS) NAIVE
Dataset
australian electricity demand 15.30 22.07 22.18 13.70 44.22 28.75 8.76 8.35 9.22 12.77 7.74 8.07 8.68 17.27 22.07
bitcoin 22.70 30.31 40.36 20.12 20.65 31.11 21.48 29.78 21.00 21.16 31.95 22.16 23.32 42.00 192.46
pedestrian counts 53.02 121.39 122.08 119.76 148.48 138.58 40.29 45.54 39.30 36.10 54.15 34.22 36.16 50.73 121.35
weather 34.05 50.85 56.19 58.06 51.47 57.98 106.01 59.12 38.17 36.46 50.87 40.13 35.35 50.85 36.81
nn5 daily 47.43 35.38 21.93 21.11 21.49 25.91 30.20 24.04 23.30 23.75 28.47 22.65 23.18 23.46 48.09
nn5 weekly 12.28 12.24 11.96 11.62 12.29 11.83 11.45 11.67 11.49 11.52 10.93 14.95 14.83 11.01 13.26
tourism yearly 48.28 34.10 31.93 33.94 36.52 33.39 46.92 31.54 33.73 34.06 30.24 28.80 34.66 30.14 42.14
tourism quarterly 25.25 27.41 15.37 17.16 15.07 16.58 15.86 16.53 16.20 15.29 14.45 15.56 16.97 19.92 31.68
tourism monthly 33.53 36.39 19.89 21.20 19.02 19.73 21.11 21.10 20.11 18.35 20.42 18.92 19.74 22.77 40.40
cif 2016 18.63 14.94 13.04 12.19 12.18 11.69 12.32 14.86 12.31 13.54 11.71 18.82 12.55 20.00 15.41
covid deaths 20.65 15.35 15.57 8.71 8.64 9.26 18.34 15.40 18.58 34.18 32.36 14.50 40.71 15.89 134.79
fred md 9.75 8.72 9.72 7.97 8.40 7.98 30.77 9.16 8.99 8.33 8.32 9.08 11.26 8.47 8.76
traffic hourly 11.70 8.73 8.73 12.58 9.84 11.72 5.97 7.94 4.30 4.16 5.16 5.22 4.10 4.76 8.73
traffic weekly 12.85 12.40 12.48 12.80 12.63 13.45 12.46 12.90 12.64 13.13 12.31 13.21 15.18 12.48 12.98
saugeenday 58.53 35.99 35.97 37.34 67.50 37.55 45.32 35.55 39.32 40.22 56.03 37.02 56.62 43.31 35.99
us births 4.48 11.77 5.82 3.81 4.05 5.17 5.75 4.23 5.55 4.13 4.17 4.88 4.36 3.94 11.35
hospital 21.68 17.94 17.27 17.55 17.46 17.79 17.56 18.04 18.29 17.41 17.72 17.51 20.04 17.86 21.55
solar weekly 37.33 24.59 24.76 19.05 22.93 17.87 21.65 29.35 21.52 15.00 24.05 32.50 12.26 24.20 32.80
In Table 2 we present the actual MAE numbers that are behind the main Figure 1(a).
It should be noted that though the scaled MAE metric accounts for scale variation across datasets, it does not account for within-dataset scale variations. Indeed, in datasets like cif 2016 and covid deaths, there are time-series with different order of magnitudes. The mean MAE for these datasets does not reflect average performance but only the performance of the higher scale time-series. Therefore in Table 3 we also present the normalized msMAPE metrics for TimesFM and all baselines.
A.2.3 Informer
@@
llmtime(ZS) PatchTST TimesFM(ZS) FEDFormer AutoFormer Informer
@@
ETTh1 (Horizon=96) 0.42 0.41 0.37 0.58 0.55 0.76
ETTh1 (Horizon=192) 0.50 0.49 0.49 0.64 0.64 0.78
ETTh2 (Horizon=96) 0.33 0.28 0.28 0.67 0.65 1.94
ETTh2 (Horizon=192) 0.70 0.68 0.58 0.82 0.82 2.02
ETTm1 (Horizon=96) 0.37 0.33 0.25 0.41 0.54 0.71
ETTm1 (Horizon=192) 0.71 0.31 0.24 0.49 0.46 0.68
ETTm2 (Horizon=96) 0.29 0.23 0.28 0.36 0.29 0.48
ETTm2 (Horizon=192) 0.31 0.25 0.24 0.25 0.30 0.51
A.3 More Details on Models
@@
TimesFM. For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280. We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $45e-$ 4.
@@
A.4 Synthetic Data
We create the synthetic data to reflect common time series patterns using traditional statistical models. We start with four simple times series patterns:

### v4

@@ Appendix A Appendix
A.1 Limitations and Future Work
Our work shows that we can train a 200M parameter pretrained forecasting model that has impressive zero-shot performance on a variety of real world forecasting benchmarks with different context and horizon lengths. In this section we would like to discuss limitations and future work.
Prompt Tuning. In LLMs it is well known that prompt tuning techniques like chain-of-thought [38] can drastically improve performance in cases where the model is inaccurate with simple prompts. Such techniques are less clear for time-series foundation model. We can tune simple hyper-parameters like context length as the moment. However, with probabilistic forecasting we might be able to output different statistics as well as come up with techniques that align more with user’s expectations while not decreasing likelihood.
Probabilistic Forecasting. It should be straightforward to train with probabilistic loss functions in our framework as detailed in the "Loss Function" part of Section 4. However, being one of the first works of building a single foundation model for forecasting, this was not our main focus and is left to future explorations. Note that as mentioned before we plan to release our model weights and after that such loss functions [30, 1] can be added during finetuning.
Covariate handling. Currently the model is not pretrained with covariates as one of the key challenges is finding large volumes of pretrained data with meaningful covariates (apart from date features). We also need methods to have a joint universal representation of covariates. Currently there are two simple techniques we can think of for handling covariates (i) In a zero-shot setting at inference time we can predict in-context and linearly regress the residual on covaraites. Then our model + the residual model can be used for forecasting in the horizon. (ii) during finetuning it is straightforward to handle covariates by adding them as inputs to the input and output residual blocks. Categorical variables can be added as embeddings.
More finetuning studies. We perform a fintuning study in Appendix A.3 following a prior work. However, a more in depth study that involves finetuning in the presence of covariates would be beneficial. This being one of the first works of building a single foundation model for forecasting, this was not our main focus and is left to future explorations. Ideas in recent work such as [4] could be useful in this regard.
Other architectures. Given the cost of training foundation models we did not perform much hyper-parameter tuning in our pretraining, while following some well established best practices for training transformers. In a similar vein, it would also be interesting to try out alternatives like the exciting directions of all MLP structures like [4] or efficient linear state space models like Mamba [10] (and references there in).
Interpretability. Deep foundation models trained on a huge corpuses of data could be inherently less interpretable compared to statistical methods like ARIMA, ETS [3]. In this regard methods like LOCO, SHAP (see [34] and references there in) could be used to some extent to attribute feature importances to different lags in the context supplied to the model. However, this does not solve the problem to a full extent and one of the best things to do would be to open source a version of the model with a proper model card. [23].
A.2 Metrics
@@
A.3 Finetuning study on ETT
In this section, we test whether TimesFM can be fintuned on a small fraction of a dataset to provide even better performance. We follow the same protocol as in GPT4TS [42] (see Table 13 in their paper). [42] finetune GPT2 input and output blocks on long-term forecasting benchmarks on 10% of the of the original datasets and compare it against models trained from scratch on the same data. Then the models are evaluated on the original test set task of [44]. We also tune the input and output residual blocks on 10% of the training set and present the results in Table 2. We can see that our model performs the best by a large margin. In ETTh1, ETTh2, ETTm1 our finetuned model is better than 18%, 3% and 12% better than GPT4TS, respectively. In fact we can see our 10% finetuned model’s performances are comparable or better than that of most baselines trained on the whole training dataset as reported in Table 14 of  [42]. This shows that the inductive biases encoded in our model weights by finetuning on a large time-series corpus are better for downstream forecasting task than an off the shelf language model like GPT2, even though our model is orders of magnitude smaller.
Table 2: MAE of different methods on ETT datasets. All methods use 10% of the original training set for training or finetuning. The baseline numbers are from Table 13 in [44].
Dataset TimesFM(FT) GPT4TS(FT) DLinear PatchTST TimeNet FEDFormer Autoformer
ETTh1 96 0.398 0.456 0.495 0.485 0.628 0.499 0.552
192 0.424 0.516 0.538 0.524 0.593 0.555 0.598
336 0.436 0.535 0.622 0.550 0.648 0.574 0.619
720 0.445 0.591 0.743 0.610 0.641 0.614 0.616
Avg 0.426 0.525 0.600 0.542 0.628 0.561 0.596
ETTh2 96 0.356 0.374 0.411 0.389 0.409 0.416 0.451
192 0.400 0.411 0.519 0.414 0.467 0.474 0.477
336 0.428 0.433 0.572 0.441 0.494 0.501 0.543
720 0.457 0.464 0.648 0.480 0.491 0.509 0.523
Avg 0.410 0.421 0.538 0.431 0.465 0.475 0.499
ETTm1 96 0.345 0.404 0.392 0.419 0.501 0.518 0.614
192 0.374 0.423 0.412 0.434 0.528 0.546 0.592
336 0.397 0.439 0.434 0.454 0.568 0.775 0.677
720 0.436 0.498 0.477 0.556 0.549 0.579 0.630
Avg 0.388 0.441 0.429 0.466 0.537 0.605 0.628
ETTm2 96 0.263 0.269 0.303 0.274 0.285 0.399 0.454
192 0.309 0.309 0.345 0.317 0.323 0.379 0.691
336 0.349 0.346 0.385 0.353 0.353 0.559 1.407
720 0.415 0.417 0.440 0.427 0.449 0.614 1.166
Avg 0.334 0.335 0.368 0.343 0.353 0.488 0.930
A.4 Pretraining PatchTST
Since TimesFM applies a similar patching strategy as PatchTST [24], for an ablation study we use the same data loader and pretrain a PatchTST model of 200M parameters to the same number of FLOPS as the final 200M TimesFM model. We denote it as PatchTST(ZS). The two models share the same hyperparameters of the transformer stack. For PatchTST(ZS) we use the same input patch length = 32, and a stride of length half of input patch size (i.e. stride = 16) as done in the original PatchTST paper.
We report the detailed results on Monash and ETT in Appendix A.5.2 and A.5.3. It can be seen that the results are not that good for PatchTST(ZS) on Monash. This is expected since our pretrain data loader will predominantly have context lengths of 512 instead of shorter context lengths as in Monash. Moreover the PatchTST model does fewer iterations at the same number of FLOPS. On ETT datasets, the PatchTST(ZS) model is performs similarly to TimesFM(ZS) and PatchTST. This is also expected since the context length for this study is indeed 512.
As PatchTST(ZS) is an encoder-decoder model, to pretrain it for zero shot forecasting one should theoretically prepare all possible context lengths and horizon lengths in the pretrain datasets. Pretraining it to its maximum performance requires much more compute and likely more careful tuning compared to pretraining TimesFM.
A.5 Additional Empirical Results
In this section, we provide more detailed tables for our zero-shot datasets and experiments described in Section 6.1. The AM based aggregated metrics are presented in Figure 4.
(a) Monash Archive [9]
(b) Darts [14]
(c) ETT (Horizons 96 and 192) [44]
Figure 4: We report average performance in three groups of datasets. In all figures, the lower the metric the better and the error bars represent one standard error. Note that among the baselines only TimesFM and llmtime are zero-shot. In (a) we report results on the Monash datasets. Since the datasets have different scales, we take the Arithmetric Mean (AM) the MAE’s scaled by the MAE of a naive baseline. We can see that TimesFM is within significance of the top model N-BEATS. In (b), we report the similarly scaled MAE on the Darts benchmarks. TimesFM is within significance of the top of method which is ARIMA in this case. Note that these datasets have one time-series each and therefore statistical methods are competitive with deep learning ones. Finally, in (c) we report the average MAE for 96 and 192 horizon prediction tasks on 4 ETT datasets i.e 8 tasks in total. TimesFM and PatchTST are the best performing models in this case.
A.5.1 Darts
We present the MAE results individually from all 8 datasets in Table 3. It can be seen that TimesFM performs well for all datasets with clear seasonal patterns. On an average we are within significant level of the best model. Note that there are only 8 time-series as a whole in Darts and theerfore these evaluations have very wide confidence intervals.
In Figure 8 we present visual comparisons of our forecasts vs some of the baselines.
@@ A.5.1 Darts
GP ARIMA TCN N-BEATS N-HiTS llmtime(ZS) TimesFM(ZS) PatchTST NAIVE
AirPassengersDataset 34.67 24.03 54.96 97.89 59.16 34.37 62.51 44.65 81.45
AusBeerDataset 102.05 17.13 30.90 10.39 34.23 16.13 11.94 21.97 96.35
GasRateCO2Dataset 2.27 2.37 2.64 2.63 3.85 3.50 2.50 2.67 2.29
MonthlyMilkDataset 30.33 37.19 70.86 33.64 32.73 9.68 28.09 42.60 85.71
SunspotsDataset 53.74 43.56 51.82 73.15 49.93 47.34 41.40 62.33 48.24
WineDataset 4552.06 2306.70 3287.14 4562.02 3909.51 1569.32 2871.33 2498.69 4075.28
WoolyDataset 649.98 588.78 1158.79 903.01 382.09 808.73 728.92 542.28 1210.33
HeartRateDataset 5.65 5.56 5.49 6.57 6.10 6.21 5.85 6.74 5.92
Scaled MAE (Arithmetic Mean) 0.8193 0.6045 0.8427 0.9176 0.8109 0.6641 0.6829 0.7462 1.0000
Scaled MAE (Geometric Mean) 0.7509 0.5219 0.7946 0.7316 0.6936 0.4882 0.5767 0.6458 1.0000
A.5.2 Monash
@@
llmtime(ZS) SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer PatchTST(ZS) TimesFM(ZS) NAIVE
@@
australian electricity demand 459.96 659.60 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.50 231.45 382.23 448.81 659.60
bitcoin 1.75e18 5.33e18 5.33e18 9.90e17 1.10e18 3.62e18 6.66e17 1.93e18 1.45e18 1.95e18 1.06e18 2.46e18 2.61e18 1.11e18 1.3e18 7.77e17
pedestrian counts 70.20 170.87 170.94 222.38 216.50 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29 51.27 40.71 170.88
weather 2.32 2.24 2.51 2.30 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03 2.07 2.07 2.36
nn5 daily 9.39 6.63 3.80 3.70 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16 3.77 3.54 8.26
nn5 weekly 15.91 15.66 15.30 14.98 15.70 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34 17.00 14.67 16.71
tourism yearly 140081.78 95579.23 90653.60 94121.08 94818.89 95033.24 82682.97 79567.22 79593.22 71471.29 70951.80 69905.47 74316.52 224411.89 109977.29 99456.05
tourism quarterly 14121.09 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67 21276.98 12102.04 15845.10
tourism monthly 4724.94 5302.10 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98 4596.21 3183.77 5636.83
cif 2016 715086.33 581875.97 714818.58 855578.40 642421.42 469059.49 563205.57 603551.30 1495923.44 3200418.00 679034.80 5998224.62 4057973.04 8374813.14 773980.44 386526.37
covid deaths 304.68 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66 348.60 209.80 353.71
fred md 2013.49 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.80 2508.40 4666.04 4965.51 947.12 2825.67
traffic hourly 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01 0.01 0.01 0.03
traffic weekly 1.17 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.20 1.42 1.23 1.12 1.19
saugeenday 28.63 21.50 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06 22.33 24.63 21.50
us births 459.43 1192.20 586.93 399.00 419.73 526.33 574.93 441.70 557.87 424.93 422.00 504.40 452.87 1193.28 437.27 1152.67
hospital 24.62 21.76 18.54 17.43 17.97 19.60 19.24 19.17 22.86 18.25 20.18 19.35 36.19 20.87 19.41 24.07
solar weekly 2049.09 1202.39 1210.83 908.65 1131.01 839.88 1044.98 1513.49 1050.84 721.59 1172.64 1996.89 576.35 1093.46 1258.27 1729.41
Scaled MAE (Arithmetic Mean) 1.0588 1.3115 1.2295 0.8824 0.9337 1.2427 1.0382 0.8924 0.8942 1.1925 0.7844 1.8119 1.4840 2.1419 0.8005 1.0000
Scaled MAE (Geometric Mean) 0.9715 1.0855 0.9371 0.7736 0.8104 0.9449 0.8218 0.7733 0.7044 0.7477 0.7005 0.9384 0.8619 1.0557 0.6846 1.0000
In Table 4 we present the actual MAE numbers that are behind the main Figure 2(a). In Figure 9, we present some examples of our zero-shot forecasts. For most datasets, we set the context window to be the maximum length of the series in the dataset capped at 512 (similar to statistcal models used in the official Monash baselines). For some datasets, we did some inference time-tuning of the context length, i.e we predict the last horizon length number of points in the training set with context lengths 32, 64 and maximum allowed and chose the best one in terms of this validation metric. This is fair as most Monash DL baselines use different context lengths for different datasets during training and our model is completely zero-shot. The max context lengths used for these datasets are (cif 2016, 32), (tourism yearly, 32), (covid deaths, 32), (bitcoin, 32), (tourism monthly, 32) and (tourism monthly, 64).
A.5.3 Informer
@@
llmtime(ZS)* PatchTST PatchTST(ZS) FEDFormer AutoFormer Informer TimesFM(ZS)
@@
ETTh1 (horizon=96) 0.42 0.41 0.39 0.58 0.55 0.76 0.45
ETTh1 (horizon=192) 0.50 0.49 0.50 0.64 0.64 0.78 0.53
ETTh2 (horizon=96) 0.33 0.28 0.37 0.67 0.65 1.94 0.35
ETTh2 (horizon=192) 0.70 0.68 0.59 0.82 0.82 2.02 0.62
ETTm1 (horizon=96) 0.37 0.33 0.24 0.41 0.54 0.71 0.19
ETTm1 (horizon=192) 0.71 0.31 0.26 0.49 0.46 0.68 0.26
ETTm2 (horizon=96) 0.29 0.23 0.22 0.36 0.29 0.48 0.24
ETTm2 (horizon=192) 0.31 0.25 0.22 0.25 0.30 0.51 0.27
Avg 0.45 0.37 0.35 0.53 0.53 0.99 0.36
A.6 More Details on Models
@@
TimesFM. For our main 200M model we use 16 attention heads, 20 layers, a input patch length of 32 and output patch length of 128. The model dimension is set to 1280. We train with layer norm and a cosine decay learning rate schedule with peak learning rate of $45e-$ 4. The hyper-parameters of TimesFM for various sizes are provided in Table 6. Note that the settings are for the base models and not ablation models. The hidden dims of both the residual block and the FFN in the transformer layers are set as the same as model dimensions. We keep layer norm in transformer layers but not in the residual blocks.
Table 6: Hyper-parameters for TimesFM
num_layers model_dims output_patch_len input_patch_len num_heads dropout
Size
200M 20 1280 128 32 16 0.2
70M 10 1024 128 32 16 0.2
17M 10 512 128 32 16 0.2
@@
A.7 Date Features
As we mentioned earlier, since we are building a single pre-trained model, we cannot have dataset specific dynamic or static covariates during training time. However, the datetime column is ubiquitous in all time-series data, so we can technically have date derived features like day of the week, month of the year etc processed into a vector at each time-point $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{r}$.
If so, the learning task can be rewritten as
$f:\mathopen{}\mathclose{{}\left(\mathbf{y}_{1:L},\mathbf{x}_{1:L+H}}\right)\longrightarrow\hat{\mathbf{y}}_{L+1:L+H}.$
There are many options to incorporate these features into the model, one being to directly concatenate them after the time-points in each patch. For this paper we decide to focus on the univariate time-series input, and will investigate this enhancement in the future.
A.8 Synthetic Data
We create the synthetic data to reflect common time-series patterns using traditional statistical models. We start with four simple times series patterns: